OpenStudy (anonymous):

What is 18 ÷ 2/3 ? A. B. 3 C. 12 D. 27

OpenStudy (anonymous):

@Rachella

OpenStudy (anonymous):

its divided by 2/3

OpenStudy (anonymous):

:) thanks i still have one more so ill close this one n tag u in it :)

OpenStudy (anonymous):

In order to solve this problem, you need to convert 18 into a fraction by simply adding a 1 as the denominator. Like this: \[\frac{ 18 }{ 1 }\div \frac{ 2 }{ 3 }\]This is a divsion problem, so we need to flip the second fraction. Then we need to multiply the numerator and denominator of the fractions to get the quotient. When we divide fractions, we have to switch the numerator and denominator of the second fraction:\[\frac{ 2 }{ 3 } \rightarrow \frac{ 3 }{ 2 }\]Now, our second fraction is: \[\frac{ 3 }{ 2 }\]Multiply the numerators together, and multiply the denominators together:\[\frac{ 18 }{ 1 }\times \frac{ 3 }{ 2 }= \frac{ 52 }{ 2 }\]The fraction 54/2 is not in lowest terms, so we need to reduce it. List the factors of the numerator and denominator: \[Factors~Of~54: 1,2,3,6,9,18,27,54\]\[Factors~Of~2: 1,2\]2 is the largest factor of both 54 and 2, so we need to divide both the numerator and denominator by 2. Numerator: 54 ÷ 2 = 27. Denominator: 2 ÷ 2 = 1. Write the new fraction:\[\frac{ 27 }{ 1 }\]Which is an improper fraction so when converted to a number, the anwser will be 27.